A rectangular blanket is 20 inches longer than it is wide. If the area of the blanket is 1,125 square inches, find the length and the width. Show all.
Answer by mananth(16949)

let the width be x
length = x+20
x(x+20) = 1125
x^2+20x-1125=0
x^2+45x-25x=1125
x(x+45)-25(x+45)=0
(x+45)(x-25)=0
x=-45 or x=25

So width = 25
and length = 20+25 = 45